OpenStudy (anonymous):

Can anyone show me how to solve for x in this problem? x+2sqrt(x) = 8

OpenStudy (anonymous):

Square every value to get rid of the radical and then solve the quadratic equation.

OpenStudy (anonymous):

\[x^2+(2\sqrt{x})^2=8^2\]

OpenStudy (lgbasallote):

note \[(x + 2\sqrt x)^2 \ne (x)^2 + (2\sqrt x)^2\] remember the FOIL ;)

OpenStudy (anonymous):

so you have \[x^2+ 4x=64\]

OpenStudy (anonymous):

thanks